Let (C*, • ) denote the group of non-zero
complex numbers and
S= {z belongs to C*| IZI = 1 }. Show that
C* / S isomorphic to R+ where (R+, • ) is the group of
positive real numbers.
1
Expert's answer
2020-12-11T12:45:24-0500

Let's consider

"\phi : \mathbb{C}^* \to \mathbb{R}^+"

"z\mapsto|z|"

This function is a surjective homomorphism.

For every "x\in\mathbb{R}^+" , we can view this number as an element of "\mathbb{C}^*" , "x\in\mathbb{C}^*, |x|=x" and thus it is surjective.

It is a homomorphism by the properties of complex numbers: "\phi(z_1z_2) = |z_1z_2|=|z_1||z_2|=\phi(z_1)\phi(z_2)"

We can also see that "Ker(\phi) = \{z\in\mathbb{C}^* : \phi(z)=|z|=1 \}=S" . Thus by the isomorphism theorem "\bar{\phi} : \mathbb{C}^*/S \to \mathbb{R}^+" is an isomorphism and so "(\mathbb{C}^*/S,\cdot) \overset{\phi}{\simeq} (\mathbb{R}^+,\cdot)" .